OpenStudy (anonymous):

Will fan and medal Solve for x x-8 / 2 = -3 @sammixboo or someone

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

thanks can you help with another

OpenStudy (anonymous):

Wait thats wrong

OpenStudy (anonymous):

A. 2 B. –9 C. –14 D.

OpenStudy (anonymous):

x-8/2=-3 right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

2, -9, -14, 7

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

working it out..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x-8/2=-3 x-4=-3 x-4+3=0 x=-3+4 x=1

OpenStudy (anonymous):